# From strong to weak correlations in breathing-mode kagome van der Waals materials: Nb3(F,Cl,Br,I)8 as a robust and versatile platform for many-body engineering

### v1

@@ VI Appendix
.1 Downfolded Model Parameters
In Table 5 we list a sub-set of the ab initio model parameters. The full parameter sets are available at TBA. We note that the Coulomb interaction has a significant long-range tail in all compounds in both the in- and out-of-plane directions.
Table 5: Molecular orbital Hubbard model parameters (hopping $t_{ij}$ and density-density Coulomb matrix elements $U_{iijj}$) in units of meV for bulk Nb3X8 in the LT structure. The lower index “0” denotes on-site values, while the upper index “(s)” and “(w)” refer to out-of-plane neighbors ($\perp$) with strong and weak bonding, respectively. The upper index $n=1,2,3$ represents the first, second, and third nearest neighbors in the in-plane direction ($\parallel$). Note that two values for $U_{\parallel}^{(n)}$ and $t_{\parallel}^{(n)}$ are provided for Nb3Cl8 due to slight asymmetry in the size of the trimers in neighboring layers within the R3 symmetry structure.
@@
$U_{0}$ 2367.2 1498.0/1493.9 1170.9 815.5
$U^{\text{(s)}}_{\perp}$ 653.7 409.8 331.7 274.8
$U^{\text{(w)}}_{\perp}$ 550.4 323.5 258.0 202.6
$U^{(1)}_{\parallel}$ 807.0 484.2/486.2 381.2 284.9
$U^{(2)}_{\parallel}$ 523.0 288.6/288.5 234.9 185.5
$U^{(3)}_{\parallel}$ 470.1 254.9/254.7 183.2 165.8
$t_{0}$ 3319.8 4166.8/4165.5 4833.0 5471.8
$t^{(\text{s})}_{\perp}$13.2 -127.0 -162.3 -207.5
$t^{(\text{w})}_{\perp}$8.8 -3.1 -19.7 -23.9
$t^{(1)}_{\parallel}$ 23.7 15.1/26.2 7.1 -10.4
$t^{(2)}_{\parallel}$2.6 5.5/4.8 5.7 6.7
$t^{(3)}_{\parallel}$ 0.4 -4.8/-4.0 -6.5 -9.8
@@
with $c=\sqrt{4\tilde{t}_{\perp}^{2}+4U\tilde{t}_{\perp}}$. To connect this simple model to our ED calculations with the full $U_{ijkl}$, we fit the self-energy of the latter using the effective dimerization $\tilde{t}_{\perp}$ as the fit parameter. We find $\tilde{t}_{\perp}\approx 0.00,0.15,0.19$ and 0.250.250.25\,0.25eV for F, Cl, Br, and I, respectively. Setting $U$ to the cRPA downfolded values $U_{iiii}$, the analytic self-energies of the simplified Hubbard dimer fit the full ED ones very well. By further comparing the Hubbard-I spectral functions obtained with and without the non-local Coulomb interactions between the two molecular orbitals of the dimer (see Appendix .3), we find the that non-local Coulomb interactions do not affect the single particle Green’s functions at half-filling.
@@
.3 Hubbard-I Spectral Functions with only on-site Coulomb Interaction
Figure 13: (a-d) Cluster DMFT spectral functions within Hubbard-I with only on-site coulomb interactions $U_{0}$ (but inter-cluster hybridization). For $N_{e}=2$ (top row) and $N_{e}=3$ (bottom row). (e) Spectral functions of the local Green’s function.
@@
For the undoped case ($N_{e}=2$), we find that the spectral functions do not change and the fundamental gaps are unaffected. This is different in the doped case ($N_{e}=3$). Here, removing the interlayer dimer density-density Coulomb interaction changes the spectral functions quantitatively in the case of Cl, Br, and I, and even qualitatively in the case of F. In the former cases, all gaps between sets of bands are modified when we neglect non-local Coulomb interactions. However, the qualitative properties are the same. Nb3Cl8 and Nb3Br8 are still charge-transfer insulators and Nb3I8 is still a (nearly) conventional Mott insulator. In the case of Nb3F8* we however find a metallic solution as soon as we neglect the non-local Coulomb interaction at $N_{e}=3$.
.4 Doping Dependence of DMFT Matsubara Self-Energies
@@
.5 Graphical Solution of the Dimer Models
@@
For $N_{e}=2$ we see three qualitatively different solutions. For Nb3F8* we find (nearly) degenerate bonding and anti-bonding self-energies, both of which have a strong pole around $\omega=0$. As a consequence of this (effectively) single-pole structure, we find two solutions to Eq. (6) (around $-1.2$ and $+1.2\,$eV) yielding two features in the spectral function. This is the classic Mott-gap scenario resulting in a lower and upper Hubbard band with a gap defined by the local Coulomb interaction strength $U$. In contrast, for the other compounds, we see that the bonding and anti-bonding self-energies still exhibit single poles, which are, however, shifted with respect to each other. As as result we find in all cases four distinct solutions, with two main spectral features around the gap and two accompanying side features. Upon close inspection we see that the upmost occupied state is given by a solution resulting from the bonding channel, and the lowest unoccupied one by a solution resulting from the anti-bonding channel. The fundamental gap is thus always opened between the bonding and anti-boding channel. But, due to the large $U$ in the case of Nb3Cl8 and Nb3Br8 the corresponding $\mathrm{Re}\Sigma(\omega)$ at the solutions $\omega=\varepsilon^{(\pm)}$ still have a finite slope. The fundamental gaps are thus significantly affected by the Coulomb interaction and especially the retardation effects imprinted into $\mathrm{Re}\Sigma(\omega)$. At a first glance, the situation seems to be similar for Nb3I8, but upon close inspection, we see that the gap is formed between $\omega=\varepsilon^{(\pm)}$ where the self-energies are nearly constant and very similar to their static values $\mathrm{Re}\Sigma(\omega=0)\approx\frac{U^{2}}{\mp 12\tilde{t}_{\perp}}$ (for the bonding and anti-bonding orbital, respectively). For the reduced $U$ and enhanced $\tilde{t}_{\perp}$ of Nb3I8, this is dominated by $\tilde{t}_{\perp}$. The gap of Nb3I8 is thus mostly controlled by static mean-field correlations.
In the doped case with $N_{e}=3$ the same analysis shows that the fundamental gap between highest occupied and lowest unoccupied states is always opened due to significant retardation effects, but solely in the anti-bonding channel and in (close) vicinity of a weakly correlated state resulting from the bonding orbital (see Fig. 15 right column). This single orbital Mott-insulator behavior shows a gradual increase in gapsize from Nb3I8 to Nb3F*8. At Nb3Cl8 the lower Hubbard band is so close in energy to the weakly correlated bonding-band that charge transfer occurs.
@@
.6 Trilayer Nb3Br8 at half-filling
@@
In the left panel of Fig. 16 we show the Wannierized DFT band structure of a trilayer Nb3Br8 using the molecular orbital basis and find three rather flat bands: a half-occupied metallic one surrounded by two fully occupied / empty ones. The metallic flat band results from the weakly hybridized “surface” monolayer and the other two are the bonding/anti-bonding splitted bands from the strongly hybridized bilayer. This is in line with the discussed band structure in Ref. [28]. As argued above, this mean-field picture is not properly reflecting the strong Coulomb interaction effects in the system. Therfore, we used this single-particle band structure together with the (molecular orbital) local Coulomb interaction matrix elements as obtained from our cRPA calculations (for the bulk material) to perform cluster DMFT calculations within the Hubbard-I approximation. The result is shown in the right panel of Fig. 16. We see that the Coulomb interactions induce a gap, with a set of three rather flat bands on either side of the Fermi energy. Projecting this interacting spectral function to the monolayer/bilayer basis shows that the gap is formed between bilayer states. In detail, the gap is formed, like in the bulk material, between the lower Hubbard band of the bonding orbital and the upper Hubbard band of the anti-bonding orbital. We also find a lower and upper Hubbard band of the initially metallic “surface” monolayer band.
@@
The two systems (“surface monolayer” and bilayer) are almost completely disentangled as a result of the small hybridization between them, consistent with the interpretation of a vdW material. We expect this electronic disentanglement to be a valid pictures for all finite-stacks with un-paired surface layers. We further note that in finite-stack systems the screening to the local Coulomb interactions will be a function of the $z$-position in the stack. The local Coulomb interaction and with it the resulting correlated gaps could thus be different in thin and thick stacks and could vary with the layer distance to the top and bottom surface.
.7 Magnetic Order in the Surface Monolayers
While the formation of singlets in the LT bulk structures of Nb3Cl8, Nb3Br8, and Nb3I8 is expected, such states cannot be formed in their surface monolayers (or HT bulk structures), nor in the LT structures of Nb3F8. This allows us to investigate the magnetic ground state properties of the surface monolayers via a classical treatment of the effective Heisenberg Hamiltonian, see Section V.2 for the computational details. For the exchange interactions in the monolayer, we use $J_{\parallel}^{(n)}$ values obtained for the bulk, see Table 2.
Our spin dynamics simulations for all monolayers of Nb3X8 compounds indicate that the ground state exhibits spin spirals. However, due to the finite size of the structures simulated, the resulting configurations yield commensurate spin spiral ordering. To address this issue, we analytically minimize Eq. (3) with respect to the wave vector $\mathbf{q}$, assuming flat spin spirals characterized by the following parametrization:
${\bf S}_{i}({\bf q})=\Big{[}\cos({\bf q}\cdot{\bf R}_{i}),\sin({\bf q}\cdot{\bf R}_{i}),0\Big{]}\,,$
where ${\bf R}_{i}$ is a position of of any classical spin ${\bf S}_{i}$.
Fig. 17 illustrates the energy of the exchange interaction (relative to the energy of the FM state) as a function of the wave vector ${\bf q}$ and the truncation distance between interacting pairs, considering up to two and three shells of nearest neighbors. The Nb3F8* and Nb3Cl8 monolayers exhibit identical propagation directions with the same value of $|{\bf q}_{\text{min}}|$, resulting in commensurate spin spirals with a periodicity of $1.5a$ (where $a$ is the in-plane lattice parameter), equivalent to a $120^{\circ}$-AFM order. While the contribution from third nearest neighbors is negligible for these two compounds, the situation changes significantly for the Nb3Br8 and Nb3I8 monolayers. Specifically, when accounting for the third nearest neighbors in Nb3Br8 and Nb3I8, the direction of spin spiral propagation is altered, and the period of modulation increases, leading to incommensurate spin configurations.
@@
Figure 17: Energy of the exchange interaction $\Delta E({\bf q})=E({\bf q})-E_{\text{FM}}$ (with respect to the energy of the FM state) as a function of the spin-spiral ${\bf q}=(q_{x},q_{y})$ vector for Nb3X8 ML. Panels (a) and (b) illustrate the results for exchange interactions $J_{ij}^{(n)}$ truncated to two ($n=1,2$) and three ($n=1,2,3$) shells of nearest neighbors, respectively. The white hexagons represent the first Brillouin zone, while the arrows indicate the ${\bf q}_{\text{min}}$ vectors corresponding to the lowest exchange interaction energy. $\lambda_{\text{min}}=2\pi/|{\bf q}|$ represents the period of the spin-spirals and $a$ is the in-plane lattice parameter.
.8 ARPES
@@
Figs. 18 (a) and (b) show the photon energy dependent ARPES data in Nb3Br8 and Nb3I8. The 2D curvature plots [107] depicted in Figs. 18 (c) and (d) reveal two dominant features which we interpret as resulting from the LHB of the bonding orbital at $\Gamma$ and $A$ in Nb3Br8 and from the upmost valence band in Nb3I8. The significant redistribution of spectral weight in energy, as visible in Figs. 18 (a) and (b) is interpreted as the result of probing different wavevectors along $\Gamma-A$ high-symmetry direction.
Plotting the energy peak positions and relative strength of the EDC fit [cf. Fig.7 (e-f)] vs photon energy once again reveals a third feature positioned lower in energy. This band is depicted by the dark blue dots for Nb3Br8 in Fig. 18 (e) and for Nb3I8 in Fig. 18 (f). We stress that this third feature is visible in Nb3I8 only at a few photon energies and, if so, with significantly reduced weight, as indicated by the dot sizes.
The feature is remarkably ubiquitous in the Nb3Br8 fits and also visible in the 2D curvature plots of Fig. 18 (c). Furthermore the band has a strongly suppressed dispersion. Based on these features we interpret this as the LHB of the anti-bonding orbital. In agreement with our observations this LHB should have suppressed dispersion and reduced spectral weight in Nb3I8 compared to Nb3Br8 (cf. Fig.7).
Figure 18: Photon energy dependent angle-resolved photoemission spectroscopy data measured over a photon energy range of 36 eV to 75 eV at $k_{\parallel}=0$. Fig. (a) shows the raw data for Nb3Br8, and (b) for Nb3I8. Panels (c) and (d) present the 2D curvature plots corresponding to the data in (a) and (b), respectively. The peak positions obtained from energy dispersion curve (EDC) fits for the valence band (VB) of Nb3Br8 and Nb3I8 are displayed in (e) and (f), respectively. Orange and red markers indicate the VB peak positions, while dark blue shows peak position of a weak band located directly below the VB. Light blue marks an additional band visible only around 38 eV for Nb3Br8 and around 56 eV for Nb3I8. Regions marked with black stars indicate areas where the bands are not clearly visible due to the presence of core levels arising from higher-order excitation. In white/yellow/black colormaps white corresponds to the minimum of the signal and black to the maximum, in blue/red/green colormaps, blue corresponds to the minimum and green to the maximum.

### v2

@@ VI Appendix
.1 Vacuum Extrapolation of the Mono- and Bilayer cRPA Coulomb Matrix Elements
For the mono- and bilayer cRPA calculations we separate the mono-/bilayers by an effective vacuum of three times their individual monolayer heights. Since the true free-standing limit would only be achieved by further increasing this effective “vacuum height”, these cRPA calculation suffer from spurious screening from the periodic replicas of the systems in c direction. To extrapolate the Coulomb matrix elements to the true free-standing limit, we make use of our Wannier Function Continuum Electrostatic (WFCE) approach [119] as applied for monolayer Nb3Cl8 in Ref. [24].
@@
$\varepsilon_{1}(q)$ $=\frac{\varepsilon_{1}^{(0)}\left[1-\tilde{\varepsilon}_{0}^{(1)}\tilde{\varepsilon}_{0}^{(2)}e^{-2qh}\right]}{1+\left[\tilde{\varepsilon}_{0}^{(1)}+\tilde{\varepsilon}_{0}^{(2)}\right]e^{-qh}+\tilde{\varepsilon}_{0}^{(1)}\tilde{\varepsilon}_{0}^{(2)}e^{-2qh}}$ with $\tilde{\varepsilon}_{0}^{(1)}=\frac{\varepsilon_{1}^{(0)}-\varepsilon_{\text{sub}}^{\text{below}}}{\varepsilon_{1}^{(0)}+\varepsilon_{\text{sub}}^{\text{below}}},\qquad\tilde{\varepsilon}_{0}^{(2)}=\frac{\varepsilon_{1}^{(0)}-\varepsilon_{\text{sub}}^{\text{above}}}{\varepsilon_{1}^{(0)}+\varepsilon_{\text{sub}}^{\text{above}}}.$ (4)
Example results for $\varepsilon_{1}(q)$ of the inital cRPA calculations for mono- and bilayer Nb3F8* are shown in Fig. 14. For both structures we see that $\varepsilon_{q=0}>1$ which is indicative for the spurious inter-layer screening due to the finite effective vacuum height. However, we also find that we can fit the raw initial cRPA data with the analytic $\varepsilon_{1}(q)$ function from above well, finding $\varepsilon_{\text{sub}}^{\text{above}}=\varepsilon_{\text{sub}}^{\text{below}}>1$ (blue curves in Fig. 14). By setting $\varepsilon_{\text{sub}}^{\text{above}}=\varepsilon_{\text{sub}}^{\text{below}}=1$ we enforce $\varepsilon_{q=0}=1$ (orange curves in Fig. 14). This way we correct for the artificial spurious inter-layer screening. Using this corrected $\varepsilon_{q}$ and the bare $v_{q}$ we can reconstruct the vacuum corrected $U_{q}$ from which we calculate in a final step via Fourier transforms the real space $U_{R}$, which are listed in Tab. 1.
Figure 14: Leading cRPA dielectric functions of monolayer (left) and bilayer (right) Nb3F8*. Markers indicate initial ab initio calculation results. Blue lines are analytic fits to the ab initio data. Orange lines represent the “vacuum corrected” behaviour. For the monolayer $\varepsilon_{1}^{(0)}$ was assumed to be a constant, while we used a Resta-like $q$-dependent model for the bilayer.
@@
with $c=\sqrt{4\tilde{t}_{\perp}^{2}+4U\tilde{t}_{\perp}}$. To connect this simple model to our ED calculations with the full $U_{ijkl}$, we fit the self-energy of the latter using the effective dimerization $\tilde{t}_{\perp}$ as the fit parameter. We find $\tilde{t}_{\perp}\approx 0.00,0.15,0.19$ and 0.250.25\,eV for F, Cl, Br, and I, respectively. Setting $U$ to the cRPA downfolded values $U_{iiii}$, the analytic self-energies of the simplified Hubbard dimer fit the full ED ones very well. By further comparing the Hubbard-I spectral functions obtained with and without the non-local Coulomb interactions between the two molecular orbitals of the dimer (see Appendix .4), we find the that non-local Coulomb interactions do not affect the single particle Green’s functions at half-filling.
@@
.3 Single-Site vs. Cluster Dynamical Mean Field Theory
Figure 16: Bulk low-temperature Nb3Br8 in a cluster and single-site Hubbard-I calculation.
In Fig. 16 we show the differences between a cluster and a single-site DMFT calculation, which respectively include the full matrix structure of the self-energy or just its diagonal components in the molecular orbital basis, for the low-temperature bulk structure of Nb3Br8. From this it is clear that to correctly describe the bonding and anti-boding Hubbard bands, cluster DMFT in a full two-molecular-orbital basis is necessary.
.4 Hubbard-I Spectral Functions with only on-site Coulomb Interaction
Figure 17: (a-d) Cluster DMFT spectral functions within Hubbard-I with only on-site coulomb interactions $U_{0}$ (but inter-cluster hybridization). For $N_{e}=2$ (top row) and $N_{e}=3$ (bottom row). Line cuts at $\Gamma$ & AA are shown in full and dashed lines respectively.
@@
For the undoped case ($N_{e}=2$), we find that the spectral functions do not change and the fundamental gaps are unaffected. This is different in the doped case ($N_{e}=3$). Here, removing the interlayer dimer density-density Coulomb interaction changes the spectral functions quantitatively in the case of Cl, Br, and I, and even qualitatively in the case of F. In the former cases, all gaps between sets of bands are modified when we neglect non-local Coulomb interactions. However, the qualitative properties are the same. Nb3Cl8 and Nb3Br8 are still charge-transfer insulators and Nb3I8 is still a (nearly) conventional Mott insulator. In the case of Nb3F8* we however find a metallic solution as soon as we neglect the non-local Coulomb interaction at $N_{e}=3$. The Nb3F8* result at three quarter filling with only on-site Coulomb interaction should be taken with care. As there is no actual dimerization in this system on might interpret it as two separate Mott insulators at non-integer filling and thus possibly away from the regime in which the Hubbard-I approximation is applicable.
.5 Full Phonon Dispersions
Figure 18: DFPT phonon dispersions for Nb3X8 in their LT bulk, bilayer, and monolayer structures.
In Fig. 18 we show the full phonon dispersion for all four compounds in their LT bulk, bilayer, and monolayer structures as calculated within DFPT.
.6 Constrained Density Functional Perturbation Theory Calculations for Nb3F8* and Nb3Cl8
Within the cDFPT calculations we removed the two lowest bands around the Fermi level from the Sternheimer equation. For detailed comparison between the DFPT and cDFPT results, we show in Fig. 19 the phonon energies of LT bulk Nb3F8* and Nb3Cl8 at the original (coarse) $q$ points without including any LO-TO splitting or applying any sum rules. The latter results in slightly imaginary frequencies at $\Gamma$ (indicated as negative values). This comparison shows that the DFPT and cDFPT results are very similar with only a few (mostly higher energy) modes being affected. As expected, those modes which change in cDFPT are shifted to higher energies, since the energy cost of atomic displacements is higher when (some) electronic states cannot move along [121]. Overall, cDFPT effects are small in Nb3Cl8 and only mildly larger in Nb3F8*. The electron-phonon coupling to the slightly gaped low-energetic flat bands of Nb3Cl8 has thus a smaller effects to the phonons than the electron-phonon coupling to the metallic flat bands in Nb3F8*.
Figure 19: Comparison between DFPT (blue/left) and cDFPT (orange/right) phonon energies of LT bulk Nb3F8* and Nb3Cl8 at selected $q$ points. Dashed lines serve as a guide to the eye and connect DFPT with corresponding cDFPT phonon energies. No acoustic sum rules are applied.
.7 Doping Dependence of DMFT Matsubara Self-Energies
@@
.8 Graphical Solution of the Dimer Models
@@
For $N_{e}=2$ we see in the left column of Fig. 21 three qualitatively different solutions. For Nb3F8* we find (nearly) degenerate bonding and anti-bonding self-energies, both of which have a strong pole around $\omega=0$. As a consequence of this (effectively) single-pole structure, we find two solutions to Eq. (7) (around 1.2-1.2 and +1.2+1.2\,eV) yielding two features in the spectral function. This is the classic Mott-gap scenario resulting in a lower and upper Hubbard band with a gap defined by the local Coulomb interaction strength $U$. In contrast, for the other compounds, we see that the bonding and anti-bonding self-energies still exhibit single poles, which are, however, shifted with respect to each other. As as result we find in all cases four distinct solutions, with two main spectral features around the gap and two accompanying side features. Upon close inspection we see that the upmost occupied state is given by a solution resulting from the bonding channel, and the lowest unoccupied one by a solution resulting from the anti-bonding channel. The fundamental gap is thus always opened between the bonding and anti-boding channel. However, due to the large $U$ in the case of Nb3Cl8 and Nb3Br8 the corresponding $\mathrm{Re}\Sigma(\omega)$ at the solutions $\omega=\varepsilon^{(\pm)}$ still have a finite slope. The fundamental gaps are thus significantly affected by the Coulomb interaction and especially the retardation effects imprinted into $\mathrm{Re}\Sigma(\omega)$. At a first glance, the situation seems to be similar for Nb3I8, but upon close inspection, we see that the gap is formed between $\omega=\varepsilon^{(\pm)}$ where the self-energies are nearly constant and very similar to their static values $\mathrm{Re}\Sigma(\omega=0)\approx\frac{U^{2}}{\mp 12\tilde{t}_{\perp}}$ (for the bonding and anti-bonding orbital, respectively). For the reduced $U$ and enhanced $\tilde{t}_{\perp}$ of Nb3I8, this is dominated by $\tilde{t}_{\perp}$. The gap of Nb3I8 is thus mostly controlled by static mean-field correlations.
In the doped case with $N_{e}=3$ (right column of Fig. 21) the same analysis shows that the fundamental gap between highest occupied and lowest unoccupied states is always opened due to significant retardation effects, but solely in the anti-bonding channel and in (close) vicinity of a weakly correlated state resulting from the bonding orbital (see Fig. 21 right column). This single orbital Mott-insulator behavior shows a gradual increase in gapsize from Nb3I8 to Nb3F*8. In the cases of Nb3F8* and Nb3Cl8 the lower Hubbard bands are so close in energy to the weakly correlated bonding-band that charge transfer occurs, rendering these compounds at these doping levels correlated charge-transfer insulators.
.9 Trilayer Nb3Br8 at half-filling
@@
In the left panel of Fig. 22 we show the Wannierized DFT band structure of a trilayer Nb3Br8 using the molecular orbital basis and find three rather flat bands: a half-occupied metallic one surrounded by two fully occupied / empty ones. The metallic flat band results from the weakly hybridized “surface” monolayer and the other two are the bonding/anti-bonding split bands from the strongly hybridized bilayer. This is in line with the discussed band structure in Ref. [30]. As argued above, this mean-field picture is not properly reflecting the strong Coulomb interaction effects in the system. Therfore, we used this single-particle band structure together with the (molecular orbital) local Coulomb interaction matrix elements as obtained from our cRPA calculations (for the bulk material) to perform cluster DMFT calculations within the Hubbard-I approximation. The result is shown in the right panel of Fig. 22. We see that the Coulomb interactions induce a gap with a set of three rather flat bands on either side of the Fermi energy. Projecting this interacting spectral function to the monolayer/bilayer basis shows that the gap is formed between bilayer states. In detail, the gap is formed, like in the bulk material, between the lower Hubbard band of the bonding orbital and the upper Hubbard band of the anti-bonding orbital. We also find a lower and upper Hubbard band of the initially metallic “surface” monolayer band.
@@
The two systems (“surface monolayer” and bilayer) are almost completely separated as a result of the small hybridization between them, consistent with the interpretation of a vdW material. We expect this electronic disentanglement to be a valid pictures for all finite-stacks with unpaired surface layers. We further note that in finite-stack systems the screening to the local Coulomb interactions will be a function of the $z$-position in the stack. The local Coulomb interaction and with it the resulting correlated gaps could thus be different in thin and thick stacks and could vary with the layer distance to the top and bottom surface.
.10 Magnetic Order in Nb3X8 Monolayers and Bulk
Following the effective spin-lattice Hamiltonian given by Eq. 2 in the main text, our spin dynamics simulations indicate that the ground state magnetic order for all bulk and monolayer of Nb3X8 exhibits conical spin spirals with in-pane propagation vector $\bf q$}
${\bf S}_{i}^{A}({\bf q},\phi^{A},\theta)=\begin{pmatrix}\sin(\theta)\cos({\bf q}\cdot{\bf R}_{i}^{A}+\phi^{A})\\ \sin(\theta)\sin({\bf q}\cdot{\bf R}_{i}^{A}+\phi^{A})\\ \cos(\theta)\end{pmatrix}\,$, (8)
where ${\bf S}_{i}^{A}$ is a classical spin of sub-lattice $A$ at site ${\bf R}_{i}^{A}$, $\theta$ is the cone angle, and $\phi^{A}$ the phase. In the absence of a magnetic field the spin spirals transforms from conical to flat with $\theta=\pi/2$. In such a case, inserting Eq. 8 into the spin-lattice Hamiltonian given by Eq. 2 results in the following expression for the energy density of the exchange interaction as a function of $\bf q$} and $\phi$
$E_{\text{ex}}({\bf q})=-\frac{1}{2N_{A}}\sum_{AB}\sum_{i,j}^{N_{A},N_{B}}J_{ij}^{AB}\cos\bigg{(}{\bf q}\cdot{\bf R}_{ij}^{AB}+\Delta\phi^{AB}\bigg{)}\,$, (9)
where the sum over $i$ is restricted to the number of magnetic sites $N_{A}$ in sublattice $A$ of the magnetic unit cell and $j$ runs over sites in sublattice $B$. ${\bf R}_{ij}^{AB}={\bf R}_{j}^{B}-{\bf R}_{i}^{A}$ is a distance between two sites and $\phi^{AB}$ is a phase difference between spirals in sublatticies $A$ and $B$. For all Nb3X8 compounds, the obtained exchange interactions $J_{ij}$ rapidly diminish beyond the third nearest neighbors in the in-plane direction and the second nearest neighbors in the out-of-plane direction. Therefore, we truncate the sum over $j$ to include only these neighboring distances. It is important to note that monolayers have only one sub-lattice and $\Delta\phi^{AB}=0$, whereas in bulk materials there are two sub-lattices and $\Delta\phi^{AB}$ may assume a finite value, potentially leading to magnetic ordering in the $z$-direction.
.10.1 Nb3X8 Monolayers
@@
Figure 23: Energy of the exchange interaction $\Delta E({\bf q})$ as a function of the spin-spiral ${\bf q}=(q_{x},q_{y})$ vector for Nb3X8 monolayers. The top and bottom panels illustrate the results for exchange interactions $J_{ij}^{(n)}$ truncated to two ($n=1,2$) and three ($n=1,2,3$) shells of nearest neighbors, respectively. The white hexagons represent the first Brillouin zone, while the arrows indicate the ${\bf q}_{\text{min}}$ vectors corresponding to the lowest exchange interaction energy. $\lambda_{\text{min}}=2\pi/|{\bf q}|$ represents the period of the spin-spirals and $a$ is the in-plane lattice parameter.
To determine the ground state parameters of the spin spirals in monolayers, we minimize the total energy described in Eq. 9 with respect to the wave vector $\bf q$}. Corresponding exchange interaction parameters $J_{ij}$ are provided in Tab. 2. Since, there is only one sublattice in monolayers, the phase $\phi$ is irrelevant to the results. Fig. 23 illustrates the energy of the exchange interaction as a function of the wave vector $\bf q$} and the truncation distance between interacting pairs, considering up to two and three shells of nearest neighbors. The Nb3F8* and Nb3Cl8 monolayers exhibit identical propagation directions with the same value of $|{\bf q}_{\text{min}}|$, resulting in commensurate spin spirals with a periodicity of $1.5a$ (where $a$ is the in-plane lattice parameter), equivalent to a $120^{\circ}$-AFM order. While the contribution from third nearest neighbors is negligible for these two compounds, the situation changes significantly for the Nb3Br8 and Nb3I8 monolayers. Specifically, when accounting for the third nearest neighbors in Nb3Br8 and Nb3I8, the direction of spin spiral propagation is altered, and the period of modulation increases, leading to incommensurate spin configurations. The final information about the ground state spirals and corresponding energies of the exchange interaction are summarized in Tab. 5.
Table 5: Magnetic ground state order in bulk and monolayer Nb3X8 compounds. Here, $\lambda=\frac{2\pi}{\bf q}$ represents the period of the spin spirals, while $\Delta\phi$ is the phase difference between the spirals in the two sublattices (applicable only to bulk structures). The parameter $\alpha_{\bf q}$ specifies the direction of spin spiral propagation relative to the in-plane unit cell vector $\bf a$}, and $E_{\text{ex}}$ denotes the corresponding energy associated with the exchange interaction (in meV).
  Bulk LT phase Bulk HT phase ML
[…]
$\lambda$ $1.5a$ $1.66a$ $3.56a$ $3.40a$ $1.72a$ $4.00a$ $1.5a$ $1.5a$ $2.8a$ $2.34a$
$\Delta\phi$ $\pi$ $\pi$ $\pi$ $\pi$ $\pi$ $\pi$
$\alpha_{\bf q}$ 0 0 0 0 0 $30^{\circ}$ 0 0 0 0
$E_{\text{ex}}$ -3.99 -37.61 -70.38 -185.98 -3.62 -3.94 -5.06 -3.99 -0.73 -3.02
.10.2 Nb3X8 Bulk
Similar to the Nb3X8 monolayers, we investigate the magnetic properties of bulk compounds in both the low- and high-temperature phases. A notable distinction from the monolayer case is the presence of out-of-plane $J^{\text{s}}_{\perp}$ and $J^{\text{s}}_{\perp}$ exchange interactions in the bulk material, see Tab. 3. Our spin dynamics simulations reveal that the ground state exhibits a spin spiral characterized by an in-plane propagation wave vector $\bf q$}. Given this preference for in-plane propagation, we analyze the spirals in the two sublattices ($A=1,2$) within each layer, which are identified by a phase $\phi^{\text{A}}$. To determine the ground state parameters of these spirals, we minimize the total energy described in Eq. 9 with respect to the wave vector $\bf q$} and the phase difference $\phi^{\text{AB}}$. The final results are summarized in Tab. 5.
Fig. 24 illustrates the energies of the exchange interactions as a function of $q$ for all LT Nb3X8 compounds, as well as HT Nb3Cl8 and Nb3Br8. The spirals associated with these energy dispersions are characterized by minimized values of $\Delta\phi$ and $\alpha_{\bf q}$, as summarized in Tab. 5. Since $\Delta\phi=\pi$ for all cases, the magnetic order at $q=0$ is AFM in the $z$-direction. From Fig. 24 it is evident that the primary contribution to the exchange energy in LT Nb3F8 arises from in-plane spiralization. Conversely, for Nb3Br8 and Nb3I8, the dominant contribution comes from out-of-plane AFM order, while the energy gain from in-plane spiralization is relatively minor. In the case of Nb3Cl8, although the energy contribution from in-plane spiralization is more significant, it remains less than that from the AFM order in the $z$-direction. Turning to the HT phases, the energy gain in Nb3Cl8, much like that in LT Nb3F8, is primarily due to in-plane magnetic ordering; meanwhile, for HT Nb3Br8, it arises from the out-of-plane AFM order.
Figure 24: Energies of the exchange interactions as a function of $\bf q$} for (a-d) all LT Nb3X8 compounds and (e-f) HT Nb3Cl8 and Nb3Br8.
.11 ARPES
Figure 25: (a),(b),(d),(e) 2D curvature plots. (c) and (f) show the photon energy dependent ARPES data over a photon energy range of 36 eV to 75 eV at $k_{\parallel}=0$.
@@
Figs. 25 (a,b,d,e) show the 2D curvature plots corresponding to the raw ARPES data in Figs. 8 (e,f,g,h). Figs. 25 (c) and (f) show the raw ARPES data at $\overline{\Gamma}$ plotted against photon energy. The significant redistribution of spectral weight in energy, as visible in Figs. 25 (c) and (f) is interpreted as the result of probing different wavevectors along $\Gamma-A$ high-symmetry direction.